OpenStudy (anonymous):

Ramesh earns 5% on a $7,500 investment compounded annually. How long will it take for the investment to reach $50,000? Compound yearly formula: f(t)=a(1+b)^t

OpenStudy (anonymous):

Can't you just plug the numbers into the equation? It's simple. Just plug 75,00 for a, .05 for b, and 50,000 for f(t).

OpenStudy (anonymous):

Yeah idk.... Is the answer 39 years orr.... 72? or 24?

OpenStudy (tkhunny):

"plug" doesn't mean anything. Substituting the appropriate values, we see that \(50000 = 7500(1.05)^{t}\) and your task is to solve for t. Unfortunately, if this doesn't come out to be an integer, it is a little misleading. Since your "answer" has three nice throws of the darts, perhaps you could demonstrate your ability to SOLVE the given equation. Go! Hint: First, divide by 7500.

OpenStudy (anonymous):

"Plug" means to substitute. I don't know if you haven't heard that before, but that's what it means. Start with using 7500 as f(t). Then, plug in the rest of your quantities for variables. Your equation should look like this: 50,000=7500(1.05)^t. Now that you've plugged it all in, you can solve the equation better. First, divide by 7500. This makes your equation look like this: 6.67=(1.05)^t.

OpenStudy (tkhunny):

If "plug" measn to substitute, why do we need the word "plug", since there is already a word that means that? "plug" means "substitute" if you want to be perceived as a 5th grader. Use "substitute" if you wish to achieve any form of credibility.

OpenStudy (anonymous):

IS the answer about 5. Please someone answer

OpenStudy (tkhunny):

The answer is nowhere near 5. \(50000=7500(1.05)^{t}\) Solve the equation. Show your work!